OpenStudy (hockeychick23):

Suppose the probability of success in a binomial event is 0.15. What is the probability of failure?

OpenStudy (hockeychick23):

(I'm assuming its 1- 0.15 but i just wanted to make sure) Thanks!

OpenStudy (campbell_st):

the 2 probabilities must add to 1 so what is the other..?

OpenStudy (hockeychick23):

Umm I'm not sure, that was the whole question

OpenStudy (anonymous):

I need help

OpenStudy (campbell_st):

the calculation so far looks correct so 240 - 52 = 188 inches now divide by 11.4 what every the answer is round down to the nearest whole number and then ths 1st cart.. if you are keen it can be done as a term in an arithmetic series with find d 109 = 52 + (6 -1)d 57 = 5d etc then a = 52 and d = 11.4 so 240 >= 52 + (n - 1)11.4 solve for n and round down to the nearest whole number... both solutions reach the same answer.